A TV manufacturer claims that the life span of its regular TV sets is 12 years with a standard



deviation of 1.2 years. Using a random sample of their 20 TV sets, the average life span is found



to be 11.2 years. Test the hypothesis that the T sets’ life span of 12 years as claimed by the



manufacturer is true at 0.05 level.

1
Expert's answer
2022-05-06T13:00:09-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=12"

"H_a:\mu\not=12"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05,"

and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{11.2-12}{1.2/\sqrt{20}}=-2.9814"

Since it is observed that "|z| = 2.981 >1.96= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-2.9814)=0.002869,"and since "p=0.002869<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 12, at the "\alpha = 0.05" significance level.